The Population of Nowheresville increased by 4% per year for the last 3 years. If the population 3 years ago was 17,350, what is the population today? Round to the nearest whole number.
Also, the same problem, but what if the population decreased by 4% for 3 years, with the starting population of 17, 350.
Thanks in advance!!

Answer by natolino_2017(77)

First Problem.
Let Pn: the population in Nowheresville after n years.
Pn = 17,350*(1,04)^n
So P3 = 17,350*(1.04)^3 = 19,516.
Second Problem.
Let Pn: the population in Nowheresville after n years.
Pn = 17,350*(0.96)^n
So P3 = 17,350*(0.96)^3 = 15,350.
For every problem, we can use a geometric progession, as the total increasing (or decreasing) is larger (smaller) every year.
@natolino_